# Large Multi-modality Model Assisted AI-Generated Image Quality Assessment

### v1

@@ 4.1. Dataset and Evaluation Metrics
Dataset. Our model is evaluated on two AI-Generated image datasets, including AIGCQA-20k (Li et al., 2024a) and AGIQA-3k (Li et al., 2023b). Specifically, AIGCQA-20k contains 20k images, but at the time of writing, only 14k images have been published. Our experiments are conducted on these 14k images. The MOS for AIGCQA-20k images are distributed between 0-5, with higher scores indicating better image quality. Images in AIGCQA-20k are generated by 15 models, including DALLE2 (Ramesh et al., 2022), DALLE3 (Ramesh et al., 2022), Dream (dreamlike art, 2023), IF (DeepFloyd, 2023), LCM Pixart (Luo et al., 2023), LCM SD1.5 (Luo et al., 2023), LCM SDXL (Luo et al., 2023), Midjourney (Holz, 2023), Pixart (Chen et al., 2023), Playground (PlaygroundAI, 2023), SD1.4 (Rombach et al., 2022b), SD1.5 (Rombach et al., 2022b), SDXL (Rombach et al., 2022a) and SSD1B (Gupta et al., 2024). AGIQA-3k includes 2982 images, with MOS also distributed between 0-5, where higher values represent better quality. Images in AGIQA-3k are derived from six models, including GLIDE (Nichol et al., 2022), Stable Diffusion V-1.5 (Rombach et al., 2022b), Stable Diffusion XL-2.2 (Rombach et al., 2022a), Midjourney (Holz, 2023), AttnGAN (Xu et al., 2018), and DALLE2 (Ramesh et al., 2022). During training, we split the entire dataset into 70 for training, 10 for validation, and 20 for testing. To ensure the same set of images in each subset when testing across different models, we set the same random seed during the split to control variables and ensure reproducibility.
Evaluation Metric. Spearman’s Rank-Order Correlation Coefficient (SRCC), Pearson’s Linear Correlation Coefficient (PLCC), the Kullback-Leibler Correlation Coefficient (KLCC), and the Root Mean Square Error (RMSE) are selected as metrics to measure monotonicity and accuracy. SRCC, PLCC, and KLCC range from -1.0 to 1.0, with larger values indicating better results. In our experiments, we employ the sum of SRCC and PLCC as the criterion for selecting the optimal validation case, and emphasize SRCC for comparing model performance.
@@ 4.2. Implementation Details
Our method is implemented based on PyTorch, and all experiments are conducted on 4 NVIDIA 3090 GPUs. For all datasets, we opt for handcrafted feature-based BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b) and ILNIQE (Zhang et al., 2015), deep learning (DL)-based HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2018), StairIQA (Sun et al., 2023), BAID (Yi et al., 2023), and LMM-based CLIPIQA (Wang et al., 2023), CLIPIQA+ (Wang et al., 2023) and Q-Align (Wu et al., 2023a). During the training process of deep learning models, we use the Adam optimizer (Kingma and Ba, 2014) with a weight decay of 1e-5, and the initial learning rate is 1e-5. The batch size is 8 during training, validation, and testing. All DL-based models are trained for 30 epochs using MSE loss and validated after each training process. The checkpoint with the highest sum of SRCC and PLCC during validation is used for testing. Handcrafted feature-based and LMM based models are used directly without training.
@@ 4.3. Comparison with SOTA methods
Table 1 lists the results of MA-AGIQA and 12 other models on the AGIQA-3k and AIGCQA-20k dataset. It has been observed that LMM-based models significantly outperform those that rely on handcrafted features. This superior performance is attributed to LMMs being trained on extensive datasets, which provides them with a robust understanding of images and enhances their generalizability. However, trained DL-based models generally perform far better than the LMM-based models because DL-based models tend to fit the data distribution of specific tasks better, thereby resulting in improved performance. Among these twelve models, the ViT-based MANIQA outperforms the other eleven models, and our method still significantly surpasses it on the same training and testing split with large margins (+3.72 of SRCC, +1.73 of PLCC and +5.43 of KRCC in AGIQA-3k & +1.61 of SRCC, +2.02 of PLCC and +2.90 of KRCC in AIGCQA-20k). This demonstrates the superiority of integrating features extracted by LMM into traditional DNN, significantly improving the accuracy and consistency of prediction results.
@@
| StairIQA | 20k3k | 0.4335 | 0.5234 | 0.3294 | 0.8549 |
| 3k20k | 0.6495 | 0.6895 | 0.4644 | 0.5285 |
To evaluate the generalization capability of our MA-AGIQA model, we conducted cross-dataset evaluations. Table 2 shows that MA-AGIQA significantly outperforms the other two models, HyperIQA and StairIQA, which performed best on single datasets, with large margins. This superior performance can largely be attributed to the robust generalization capability of the LMM and the benefits of the MoE architecture, which excels in dynamically fusing features.
@@ 4.4. Ablation Study
Necessity of Fine-grained Semantic Features. To assess the benefits of integrating features extracted by mPLUG-Owl2 (Ye et al., 2023) into MANIQA (Yang et al., 2022), we carried out comprehensive ablation studies on each component and their various combinations, as detailed in Tables 3 and 4. Our findings indicate that using either the features extracted by the LMM alone or solely relying on a traditional network does not yield the best outcomes. In contrast, integrating one fine-grained semantic feature with the original MANIQA network can enhance the network’s performance. However, the optimal results were achieved by combining two features extracted by the LMM with MANIQA, which led to significant improvements on the AGIQA-3k dataset (increases of 1.57, 0.83, and 2.56 in SRCC, PLCC, and KRCC, respectively) and on the AIGCQA-20k dataset (enhancements of 2.72, 1.94, and 4.35).
@@
Contribution of MoE. Table 5 demonstrates that incorporating the MoE structure, rather than simply concatenating three vectors, does indeed improve network performance, albeit marginally. Specifically, on the AGIQA-3k dataset, we observed increases of 0.20, 0.17, and 0.16 in SRCC, PLCC, and KRCC, respectively. For the AIGCQA-20k dataset, the improvements were 0.67, 0.95, and 1.37. The gains, although seemingly modest, highlight the potential of MoE structure in complex systems where integrating diverse expertise can yield better decision-making and predictive outcomes.
@@ 4.5. Visualization
To vividly demonstrate the efficacy of the MA-AGIQA framework, we selected 300 images from the AIGCQA-20k and AGIQA-3k datasets where MANIQA had the poorest performance. These images primarily exhibit issues in semantic content. We computed the absolute values of the differences between the model scores and the image ground truth, and illustrated these differences in Figure 6, using 0.1 as the bin size for plotting the quality score distribution. The results clearly show that our MA-AGIQA model are more closely aligned with human perception, with a noticeable shift in the difference distribution toward zero and a marked reduction in peak values.
Figure 7 presents a collection of images where the assessments from the MANIQA model were mostly off the mark. Scores assigned by MANIQA alongside those given by the proposed MA-AGIQA model and the ground truth are listed, which reveal that the MA-AGIQA model markedly enhances alignment with the ground truth in contrast to MANIQA. For instance, in the first image of the top row, MANIQA’s score is 3.50, which diverging substantially from the ground truth score of 1.50. However, MA-AGIQA’s score is 2.98, demonstrating a much closer approximation to the ground truth. This pattern is consistent across the images shown, with MA-AGIQA consistently producing scores that are closer to the ground truth, reflecting a more accurate assessment of image quality.
@@ 5. Conclusion
To mitigate the shortcomings of traditional DNNs in capturing semantic content in AGIs, this study explored the integration of LMMs with traditional DNNs and introduced the MA-AGIQA network. Leveraging mPLUG-Owl2 (Ye et al., 2023), our network efficiently extracts semantic features to enhance MANIQA (Yang et al., 2022) for quality assessment. The MA-AGIQA network’s ability to dynamically integrate fine-grained semantic features with quality-aware features enables it to effectively handle the varied quality aspects of AGIs. Experiment results across two prominent AGIs datasets confirm our model’s superior performance. Through thorough ablation studies, the indispensable role of each component within our framework has been validated. This research aspires to catalyze further exploration into the fusion of LMMs within AI-generated content quality assessment and envisions broader application potentials for such methodology.

### v2

@@ 4.1. Dataset and Evaluation Metrics
Dataset. Our model is evaluated on two AI-Generated image datasets (Li et al., 2023b, 2024a) and two traditional IQA datasets (Hosu et al., 2020; Fang et al., 2020). Note that AIGCQA-20k contains 20k images, but at the time of writing, only 14k images have been published and our experiments are conducted on these 14k images. Images in AIGCQA-20k are generated by 15 models, including DALLE2 (Ramesh et al., 2022), DALLE3 (Ramesh et al., 2022), Dream (dreamlike art, 2023), IF (DeepFloyd, 2023), LCM Pixart (Luo et al., 2023), LCM SD1.5 (Luo et al., 2023), LCM SDXL (Luo et al., 2023), Midjourney (Holz, 2023), Pixart (Chen et al., 2023), Playground (PlaygroundAI, 2023), SD1.4 (Rombach et al., 2022b), SD1.5 (Rombach et al., 2022b), SDXL (Rombach et al., 2022a) and SSD1B (Gupta et al., 2024). AGIQA-3k includes 2982 images. Images in AGIQA-3k are derived from six models, including GLIDE (Nichol et al., 2022), SD1.5, SDXL-2.2, Midjourney, AttnGAN (Xu et al., 2018), and DALLE2. During training, we split the entire dataset into 70 for training, 10 for validation, and 20 for testing. To ensure the same set of images in each subset when testing across different models, we set the same random seed during the split to ensure reproducibility.
Evaluation Metric. Spearman’s Rank-Order Correlation Coefficient (SRCC), Pearson’s Linear Correlation Coefficient (PLCC), the Kullback-Leibler Correlation Coefficient (KLCC), and the Root Mean Square Error (RMSE) are selected as metrics to measure monotonicity and accuracy. SRCC, PLCC, and KLCC range from 0 to 1.0, with larger values indicating better results. In our experiments, we employ the sum of SRCC and PLCC as the criterion for selecting the optimal validation case, and emphasize SRCC for comparing model performance.
@@ 4.2. Implementation Details
Our method is implemented based on PyTorch, and all experiments are conducted on 4 NVIDIA 3090 GPUs. For all datasets, we opt for handcrafted feature-based BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b) and ILNIQE (Zhang et al., 2015), DNN-based HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2018), StairIQA (Sun et al., 2023), BAID (Yi et al., 2023), and LMM-based CLIPIQA (Wang et al., 2023), CLIPIQA+ (Wang et al., 2023) and Q-Align (Wu et al., 2023a). During the training process, all DNN-based models are trained for 30 epochs using MSE loss and validated after each training process. Handcrafted feature-based and LMM based models are used directly without training.
@@ 4.3. Comparison with SOTA methods
Table 1 lists the results of MA-AGIQA and 12 other models. On AGIs dataset, it has been observed that LMM-based models significantly outperform those that rely on handcrafted features. This superior performance is attributed to LMMs being trained on extensive datasets, which provides them with a robust understanding of images and enhances their generalizability. However, trained DNN-based models generally perform far better than the LMM-based models because DNN-based models tend to fit the data distribution of specific tasks better, thereby resulting in improved performance. Among these twelve models, the ViT-based MANIQA outperforms the other eleven models, and our method still significantly surpasses it on the same training and testing split with large margins. Despite being specifically designed for AGIs, MA-AGIQA outperforms all the compared methods on both NSIs datasets. This suggests that semantic information is also crucial for evaluating the quality of NSIs. It is worth noting that semantic information plays a more significant role in AGI quality assessment than in NSI quality assessment, as evidenced by the larger performance gap on AGIs datasets.
@@
MANIQA | 20k3k | 0.7653 | 0.7980 | 0.5681 | 0.6041 |
3k20k | 0.7301 | 0.7975 | 0.5372 | 0.4402 | |
MA-AGIQA | 20k3k | 0.8053 | 0.8430 | 0.6083 | 0.5399 |
3k20k | 0.7722 | 0.8314 | 0.5777 | 0.4055 |
To evaluate the generalization capability of our MA-AGIQA model, we conducted cross-dataset evaluations. Table 2 shows that MA-AGIQA significantly outperforms the other two models, HyperIQA and MANIQA, which performed best on single datasets, with large margins. This superior performance can largely be attributed to the robust generalization capability of the LMM and the benefits of the MoE architecture, which excels in dynamically fusing features.
@@ 4.4. Ablation Study
Necessity of Fine-grained Semantic Features. To assess the benefits of integrating features extracted by mPLUG-Owl2 (Ye et al., 2023) into MANIQA (Yang et al., 2022), we carried out comprehensive ablation studies on each component and their various combinations, as detailed in Tables 3 and 4. Our findings indicate that using either the features extracted by the LMM alone or solely relying on a traditional network does not yield the best outcomes. In contrast, integrating one fine-grained semantic feature with the original MANIQA network can enhance the network’s performance. However, the optimal results were achieved by combining two features extracted by the LMM with MANIQA, which led to significant improvements on two AGIs datasets.
@@
Contribution of MoE. Table 5 demonstrates that incorporating the MoE structure, rather than simply concatenating three vectors, does improve network performance, albeit marginally. Specifically, on the AGIQA-3k dataset, we observed increases of 0.20, 0.17 in SRCC and PLCC, respectively. For the AIGCQA-20k dataset, the improvements were 0.67 and 0.95. The gains, although seemingly modest, highlight the potential of MoE structure in complex systems where integrating diverse expertise can yield better decision-making and predictive outcomes.
@@
4.5. Computational Costs
Although integrating LMM features with traditional DNN-based models does introduce additional cost, we emphasize that the incremental cost is minimal: First, the LMM model does not require training and the trained parameters and GFLOPs of ours is similar to those of traditional DNN-based method MANIQA (Parameters: M vs. M, FLOPs: G vs. G). Second, the inference time of extracting LMM features for an image with a resolution of only needs seconds on a NVIDIA RTX 3090 GPU.
@@ 4.6. Visualization
To vividly demonstrate the efficacy of the MA-AGIQA framework, we selected 300 images from the AIGCQA-20k and AGIQA-3k datasets where MANIQA had the poorest performance. These images primarily exhibit issues in semantic content. The absolute values of the differences between the model scores and the image ground truth are illustrated in Figure 5, using 0.1 as the bin size for plotting the quality score distribution. The results clearly show that our MA-AGIQA model are more closely aligned with human perception, with a noticeable shift in the difference distribution toward zero and a marked reduction in peak values.
Figure 6 presents a collection of images where the assessments from MANIQA were mostly off the mark. Scores assigned by MANIQA alongside those given by the proposed MA-AGIQA and the ground truth are listed, which reveal that MA-AGIQA markedly enhances alignment with the ground truth in contrast to MANIQA. For instance, in the first image of the top row, MANIQA’s score is 3.50, which diverging substantially from the ground truth score of 1.50. However, MA-AGIQA’s score is 2.98, demonstrating a much closer approximation to the ground truth. This pattern is consistent across all of the images, with MA-AGIQA consistently producing scores that are closer to the ground truth.
@@ 5. Conclusion
To mitigate the shortcomings of traditional DNNs in capturing semantic content in AGIs, this study explored the integration of LMMs with traditional DNNs and introduced the MA-AGIQA network. Leveraging mPLUG-Owl2 (Ye et al., 2023), MA-AGIQA efficiently extracts semantic features to enhance quality assessment. The MA-AGIQA network’s ability to dynamically integrate fine-grained semantic features with quality-aware features enables it to effectively handle the varied quality aspects of AGIs. Experiment results across four datasets confirm our model’s superior performance. With thorough ablation studies, the indispensable role of each component within our framework has been validated. This research aspires to catalyze further exploration into the fusion of LMMs within AI-generated content quality assessment and envisions broader application potentials for such methodology.
6. Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62225112, 62101326, 62301316, and 62271312, the China Postdoctoral Science Foundation under Grants 2023TQ0212 and 2023M742298, and the Postdoctoral Fellowship Program of CPSF under Grant GZC20231618.